Status Report

NOAA GOES SATOPS Morning Report 4 October 2000

By SpaceRef Editor
October 4, 2000
Filed under

SATOPS MORNING REPORT – October 4, 2000

POES

No change in the status of the POES spacecraft.

NOAA-16 operations nominal.  AMSU OV tests continue.

POES operations were nominal over the past 24 hours.

DMSP

No change in the status of the DMSP spacecraft.

DMSP operations were nominal over the past 24 hours.

GOES

No change in the status of the GOES spacecraft.

GOES-8 (East) Operations:

The GOES-8 fall eclipse season occurs daily between 0414-0526Z and continues
through October 15th.  GOES-8 images cancelled for Eclipse/KOZ – 0345,
0401, 0409, 0415, 0431, 0439, 0445, 0501, 0509, 0515, 0531, 0539, 0545Z
and 0615Z.  GOES-8 soundings cancelled for Eclipse/KOZ – 0346, 0420,
0446, 0520, 0546 and 0620Z.  Full disk coverage during GOES-10 KOZ:
0815Z – 1044Z.

After the reconfiguration to the 2209 MHz telemetry stream (for anticipated
solar RFI), the 1646Z sounding was lost after a schedule suspend due to
intermittent telemetry.  The received 2209 MHz signal was reported
as weak at the CDA.  The 1694 MHz stream was reconfigured at 1647Z
without event.

GOES-10 (West) Operations:

The GOES-10 fall eclipse season occurs daily between 0816-0930Z and
continues through October 16th.  GOES-10 images cancelled for KOZ/Eclipse:
0800, 0815, 0822, 0830, 0845, 0851, 0900, 0930, 1021 and 1030Z.  GOES-10
soundings cancelled for KOZ/Eclipse: 0801, 0824, 0901, 0924 and 1024Z. 
Full disk coverage during GOES-8 eclipse 0430Z – 0629Z.

At 1748Z on Tuesday, October 3, a GOES-10 frame break was observed and
attributed to the recycling of SPS 4 at the CDA.

Between 1752Z on Tuesday, October 3 and 0100Z on Wednesday, October
4, multiple GOES-10 frame breaks were observed and attributed to high BER
at the CDA (unknown reason).  Actual frame breaks were recorded at
1752Z, 1800Z, 1830Z, 1845Z, 2353Z, 0009Z, 0038Z and 0100Z.

At 0122Z on Wednesday, October 4, a GOES-10 frame break was observed
and attributed to loss of signal synch.

At 0239Z and 0324Z on Wednesday, October 4, GOES-10 frame breaks were
observed and attributed to low level RFI at the CDA (unknown source).

Storage Mode Spacecraft Operations:

The GOES-9 fall eclipse season occurs daily between 0612-0724Z and continues
through October 15th.  GOES-9 remains in the Z-Axis Precession (ZAP)
storage mode.

The GOES-11 fall eclipse season occurs daily between 0619-0732Z and
continues through October 16th.  GOES-11 remains in the Z-Axis Precession
(ZAP) storage mode.

GOES operations were nominal over the past 24 hours.

 

Questions or comments to twalsh@nesdis.noaa.gov

SpaceRef staff editor.